OpenStudy (anonymous):

P w l 14 2 ? 22 8 ? Solve P = 2(l + w) for l. What are the missing values in the table.

OpenStudy (anonymous):

l + w = P / 2 l = (P/2) - w missing values are 14/2 - 2 = 5 22/2 - 8 = 3